the diagonal of a square is 62 feet. What is the perimeter?
Answer by oberobic(2304)

The diagonal of a square is the hypotenuse of a triangle.
.
c^2 = a^2 + b^2
.
Since it is a square, a=b.
.
62^2 = a^2 + b^2 = a^2 + a^2
.
62^2 = 2a^2
.
3844 = 2a^2
.
2a^2 = 3844
.
a^2 = 1922
.
a = sqrt(1922)
.
a = 43.84
.
So, the perimeter = 4*43.84 = 175.35 or about 176
.
Done